OpenStudy (anonymous):

If 4^2-x=3, then 4^2x-1=? Step by step solutions pls

OpenStudy (anonymous):

4^2-x=3 16-x=3 x=13 4^2 (13)-1=x 208-1=x x=207

OpenStudy (anonymous):

\[4^{2}-x = 3\]\[16-x = 3\]\[-x = -13\]\[x=13\]Then just plug in x into the 2nd equation and solve:\[4^{2 \times 13} - 1 = y\]\[4503599627370495 = \]

OpenStudy (anonymous):

y

OpenStudy (anonymous):

oh sorry i misread the question

OpenStudy (anonymous):

Your answer seems a lot less "out there." I solved the question that was posted, but the question was probably formatted incorrectly, so you're probably right.

OpenStudy (anonymous):

is it 4^(2-x) or (4^2)-x

OpenStudy (anonymous):

uh huh thankx

OpenStudy (anonymous):

then alex is right

OpenStudy (anonymous):

if the equation was 4^(2-x)=3 then : (4^(2-x))^2=3^2 4^(4-2x)=9 4^4=9*4^2x 4^3=9*4^(2x-1) 4^(2x-1)=64/9

OpenStudy (anonymous):

why did you square the 4^(2-x) and the 3? but you got the right answer.

OpenStudy (anonymous):

If 4^2-x=3, then 4^2x-1=? Step by step solutions pls \[4^2-x=3\] \[16-x=3\] \[-x=3-16\] \[-x=-13\] \[x=13\]\[4^2x-1=16(13)-1=207\]